# The Validation Gap: A Mechanistic Analysis of How Language Models Compute Arithmetic but Fail to Validate It

### v1

@@ 4.2 Decoding the Error Detection Process
Once we obtain a soft intersection circuit $\mathcal{C}^{(\tau)}$, we can analyze its components to gain deeper insights into the model’s error detection mechanisms. For instance, $\mathcal{C}_{\text{answer}}^{(8/8)}$ contains several edges that connect attention heads in the model’s middle layers at the position of the error, as illustrated in Figure 4. To better understand the function of these attention heads, we compute their average attention scores over a set of input prompts. Figure 5 shows the average attention scores of the second attention head in layer 12 (present in Qwen-2.5-1.5B’s $\mathcal{C}_{\text{result}}^{(5/8)}$ circuit). Specifically, we visualize attention scores for four different sets of prompts: i) prompts with an error at the position of the arithmetic result, ii) prompts with an error at the position of the final numeric answer, iii) prompts without errors, and iv) prompts with a consistent error at both the arithmetic result and the final numeric answer (e.g., “…$5+8=16$. Thus, Jane has $16$ apples.”). Two notable attention patterns emerge. For prompts where an error is present either at the position of the arithmetic result (e.g., “…$5+8=16$. Thus, Jane has $13$ apples.”) or at the position of the final numeric answer (e.g., “…$5+8=13$. Thus, Jane has $16$ apples.”), we observe high attention scores between the first digit of the arithmetic result ([result-first]) [^5] Note that Qwen-2.5 uses a one-digit tokenization scheme; i.e., the number $13$ is encoded into two separate tokens. and the first digit of the final numeric answer ([answer-first]), but not for the corresponding second digits ([result-second] and [answer-second], respectively). In contrast, for prompts without errors or those with consistent errors at both positions (e.g., “…$5+8=16$. Thus, Jane has $16$ apples.”), we observe high average attention scores for both the first and second digits of the result and the final numeric answer. In essence, we observe that the attention head exhibits high average attention scores when the digits of the arithmetic result and the final numeric answer align. We refer to such attention heads as consistency heads—attention heads that assess surface-level alignment of numerical values in the arithmetic solution. Notably, we find several consistency heads across all models (see Figures 15(a) and 16(a) in the Appendix). Additionally, we observe (in)consistency heads, which display high attention scores for numerical values that misalign (Figures 13(b) to 16(b) in the Appendix).
Figure 6: Consistency head patching. Accuracy of Qwen-2.5-1.5B on (clean, corrupt) prompt pairs, where the clean prompt contains a consistent error at both error positions. The blue intervention bar represents the result after patching a set of six consistency heads (see Table 3 in the Appendix). In contrast, the orange bar indicates the accuracy after patching six randomly chosen attention heads that are not labeled as consistency heads.
@@
We hypothesize that consistency heads in the lower to middle layers of the models play an important role in the arithmetic error detection process. This hypothesis has two key implications: i) models may struggle to distinguish between samples without errors and those containing a consistent error at both the position of the arithmetic result and the final numeric answer (e.g., “…$5+8=16$. Thus, Jane has $16$ apples.”); and ii) a small subset of consistency heads can significantly influence the models’ arithmetic error detection behavior. To test our hypothesis, we first evaluate all models on 1,000 (clean, corrupt) prompt pairs for each template, where the clean prompts contain a consistent error at both error positions. Our findings reveal that all models exhibit significant difficulties in detecting these type of errors. For instance, Qwen-2.5-1.5B achieves an average accuracy of only $12.39\pm 6.00$, exhibiting a strong bias toward labeling clean prompts as “valid” (detailed results for all models can be found in Table 6 in the Appendix). Next, we analyze the impact of individual consistency heads on the model’s error detection behavior. When running the model on a prompt with a consistent error at both error positions, we patch (or substitute) the latent activations of six consistency heads (see Table 3 in the Appendix) with the corresponding activations obtained from a prompt where the error is present only at the arithmetic result (for which the model can detect the error). Figure 6 shows the resulting change in accuracy for Qwen-2.5-1.5B, demonstrating that a limited number of consistency heads can effectively control the model’s error detection behavior. Interestingly, similar results are observed for all other models (see Figure 17 in the Appendix). For details on the patching procedure and the results for other models, refer to Appendix D.4 and E.5, respectively.
@@
Figure 7: Probing. The linear probe’s accuracy across all layers of Qwen-2.5-1.5B at selected token positions. Only in higher layers, the probe is able to successfully predict the correct arithmetic result.
@@ 4.3 Dissociation of Arithmetic Validation and Computation
Our findings presented in Section 4.2 suggest that the considered models tend to rely on surface-level consistency checks rather than re-evaluating the given arithmetic equation and comparing the result with the final numeric answer. Notably, we find that all models achieve 100% accuracy in predicting the correct result of the arithmetic equation provided in each prompt. However, we hypothesize that this correctly predicted result is not used for validation. To better understand the relationship between the models’ arithmetic computation and validation procedures, we identify another set of circuits responsible for computing the correct arithmetic result at the position of the arithmetic equation (e.g., $5+8=13$), following a similar procedure as described in Section 3. For further details on the identification process, please refer to Appendix D.3. Interestingly, we find that for all models, the identified arithmetic circuits predominantly contain edges in higher layers (due to space constraints, visualizations of these circuits are provided in Figures 20 to 23 in the Appendix). This structural dissociation between the circuits responsible for arithmetic computation and those involved in validation seems to explain the models’ difficulties in detecting even simple arithmetic errors. Specifically, although the models can successfully re-compute the result of a given arithmetic equation, the final arithmetic outcome is not fully encoded when the model checks for numeric consistency in middle layers. To support this hypothesis, we train a linear probe to predict the correct arithmetic result based on the hidden states of the model’s residual stream (for training details, see Appendix D.5). Figure 7 shows the probe’s accuracy across different layers of Qwen-2.5-1.5B and selected token positions. Only in the higher layers does the model’s residual stream linearly encode information about the correct arithmetic result. Interestingly, similar patterns are observed for other models, too (see Figure 18 in the Appendix).
@@
Figure 8: Closing the validation gap. Accuracy of Qwen-2.5-1.5B on (clean, corrupt) prompt pairs where (left) the clean prompt contains a consistent error at both error positions, and (right) an error is present only at the position of the arithmetic result. The blue intervention bar denotes the result after adding the hidden representation of the residual stream in layer 22 (at [result-first]) to the residual stream of layer 1 (at [result-second]).
@@
Finally, we demonstrate that by “bridging” the gap between arithmetic computation and validation, the model’s error detection capacity for prompts with consistent errors at both error positions can be significantly enhanced. Figure 8 shows the error detection accuracy of Qwen-2.5-1.5B after we add the hidden representation of the residual stream from layer 22 at the position of the arithmetic result’s first digit to the residual stream of layer 1 at the position of the second-digit. Notably, this approach achieves an accuracy improvement of around 81%. Furthermore, accuracy on samples containing errors only at the position of the arithmetic result does not decline markedly, indicating that instead of overwriting existing information, the added information complements and enhances the model’s error detection capabilities. A similar trend for all models is presented in Figure 19 in the Appendix.
@@ 5 Related Work
Self-correction in LLMs refers to the ability of models to correct their own generated output (Kamoi et al., 2024b; Huang et al., 2024; Madaan et al., 2023). Recent studies (Tyen et al., 2024; Kamoi et al., 2024a) suggest that LLMs tend to struggle with intrinsic self-correction, especially with detecting errors in their own output (Huang et al., 2024; Tyen et al., 2024; Kamoi et al., 2024b). While most studies focus on improving the models’ ability to self-correct (Kamoi et al., 2024b; Madaan et al., 2023; Chen et al., 2024; Zhao et al., 2023; Welleck et al., 2023), our study analyzes the capacity of LLMs to detect errors from a mechanistic point of view.
@@
To the best of our knowledge, the underlying processes of arithmetic reasoning and error detection have been studied independently in LLMs so far. Several studies (Stolfo et al., 2023; Zhang et al., 2024; Nikankin et al., 2024) use causal mediation analysis (Pearl, 2001) to identify circuits that account for how LLMs process arithmetic operations. As of now, only few studies have analyzed self-correction in LLMs beyond the models’ generated output (Li et al., 2024; Liu et al., 2024b).
@@ 7 Limitations
As discussed in Section 3, our analysis is limited to four smaller-sized LLMs. Although we observe consistent patterns across various model architectures, sizes, and fine-tuning procedures (particularly within the mathematical domain), future research could investigate how these findings extend to larger models with more advanced arithmetic capabilities.
@@ Acknowledgments
We would like to thank the members of the MaiNLP lab for their valuable feedback, with special thanks to Michael Hedderich, Robert Litschko, Diego Frassinelli, Rob van der Goot, Siyao Peng, Yang Janet Liu, Xinpeng Wang, Verena Blaschke, Raoyuan Zhao, Elena Senger, Felicia Körner, Soh-Eun Shim, Andreas Säuberli, Florian Eichin, Shijia Zhou, and Beiduo Chen. Additionally, we appreciate the insightful comments and suggestions provided by the anonymous reviewers. RB is grateful to Amazon Alexa for the generous donation used to support this work. Finally, we acknowledge the support for BP through the ERC Consolidator Grant DIALECT 101043235.
@@ B.3 Circuit Identification Process
Appendix C Dataset
C.1 Templates
We generate our dataset of clean and corrupt prompts based on the templates in Table 7. These templates have a set of variables, namely [instruction], [person], [object], [pronoun], [num1], [num2], [num3], each of which can be assigned different values. Table 8 lists all possible values. For the numerical variables, we use single-digit numerical values ([num1] and [num2]) that add to a two-digit arithmetic result ([num3]). To ensure that each variable assignment occupies the same position in the token sequence within a template, we retain only variables that are tokenized as a single token for each model. For the [instruction] variable, we include only instructions that share the same number of tokens. Finally, for the [correct_pair] variable, we select assignments where labels are tokenized as a single token across models.
C.2 Aligning Token Positions Across Templates
Appendix D Experiment Details

### v2

@@ 4.2 Decoding the Error Detection Process
Once we obtain a soft intersection circuit $\mathcal{C}^{(\tau)}$, we can analyze its components to gain deeper insights into the model’s error detection mechanisms. For instance, Figure 4 illustrates that $\mathcal{C}_{\text{answer}}^{(8/8)}$ contains several edges that connect attention heads in the model’s middle layers 8 to 15 at the position of the error ([answer-second]). To better understand the function of such attention heads, we compute their average attention scores over a set of input prompts. Figure 5 shows the average scores of the second attention head in layer 12 present in Qwen-2.5-1.5B’s $\mathcal{C}_{\text{result}}^{(5/8)}$ circuit. Specifically, we visualize attention scores for four different sets of prompts: i) prompts with an error at the position of the arithmetic result, ii) prompts with an error at the position of the final numeric answer, iii) prompts without errors, and iv) prompts with a consistent error at both the arithmetic result and the final numeric answer (e.g., $5+8={\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16}$. Thus, Jane has $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16$} apples.”). Two notable attention patterns emerge. For prompts where an error is present either at the position of the arithmetic result (e.g., $5+8={\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16}$. Thus, Jane has $13$ apples.”) or at the position of the final numeric answer (e.g., “…$5+8=13$. Thus, Jane has $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16$} apples.”), we observe high attention scores between the first digit of the arithmetic result ([result-first]) and the first digit of the final numeric answer ([answer-first]), but not for the corresponding second digits ([result-second] and [answer-second], respectively). In contrast, for prompts without errors or those with consistent errors at both positions (e.g., $5+8={\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16}$. Thus, Jane has $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16$} apples.”), we observe high average attention scores for both the first and second digits of the result and the final numeric answer. In essence, we observe that the attention head exhibits high average attention scores when the digits of the arithmetic result and the final numeric answer align. We refer to such attention heads as consistency heads—attention heads that assess surface-level alignment of numerical values in the solution prompt. Notably, we find several consistency heads across all models (see Figures 16(a) and 17(a) in the Appendix). Additionally, we observe (in)consistency heads, which display high attention scores for numerical values that misalign (Figures 14(b) to 17(b) in the Appendix).
@@
We hypothesize that consistency heads in the lower to middle layers of the models play an important role in the arithmetic error detection process. This hypothesis has two major implications: i) models may struggle to distinguish between samples without errors and those containing a consistent error at both the position of the arithmetic result and the final numeric answer (e.g., $5+8={\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16}$. Thus, Jane has $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16$} apples.”); and ii) a small subset of consistency heads can significantly influence the models’ arithmetic error detection behavior. To test our hypothesis, we first evaluate models on 1,000 (clean, corrupt) prompt pairs for each template, where clean prompts contain a consistent error at both error positions. As shown in Table 2, all models exhibit significant difficulties in detecting these type of errors. For instance, Qwen-2.5-1.5B achieves an average accuracy of only $12.39\%\pm 6.00\%$, indicating a strong bias toward labeling prompts with a consistent error as “valid”. Even for Phi-3-Mini-4k-Instruct, the accuracy drops from about 89% when detecting errors at either the arithmetic result or the final numeric answer (see Table 1) to $40.98\%\pm 18.41\%$ when both positions contain a consistent error.
Next, we analyze the influence of individual consistency heads on the model’s error detection behavior through two complementary experiments. First, when Qwen-2.5-1.5B is given a prompt with a consistent error at both positions (for which it incorrectly predicts “valid”), we patch the latent activations of six consistency heads (see Table 4 in the Appendix for the exact heads) with the corresponding activations from a prompt containing a single error at the arithmetic result (for which the model correctly predicts “invalid”). If consistency heads indeed govern error detection, the model should change its initial prediction from “valid” to “invalid” under this intervention. Second, we perform the reverse: running the model on a single-error prompt (correctly predicted as “invalid”), we patch the activations of the consistency heads with those from a consistent-error prompt. In this case, we would expect the intervention to reduce the error detection accuracy.
Figure 6: Error detection accuracy of Qwen-2.5-1.5B before and after patching six consistency heads. On the left, we evaluate the model on consistent-error prompts and patch activations from a single-error run; on the right, we evaluate on single-error prompts and patch with activations from a consistent-error run. As a control, we patch six randomly chosen attention heads.
Figure 6 shows the resulting changes in detection accuracy (predicting “invalid”) for Qwen-2.5-1.5B. As expected, the first intervention $\left(\text{ConsErr}[\text{heads}:=\text{SingErr}]\right)$ reliably flips the model’s prediction from “valid” to “invalid,” demonstrating that inconsistency signals injected via these heads can causally alter the model’s output. By contrast, the reverse intervention $\left(\text{SingErr}[\text{heads}:=\text{ConsErr}]\right)$ changes the prediction from “invalid” to “valid” only in some cases. Similar results hold across all models (see Figure 18 in the Appendix). A plausible explanation might be that the model has more consistency heads than the ones we patch, which still signal an inconsistency for the given prompt. This touches upon the problem of completeness in circuit discovery, where circuits that can faithfully reproduce the model’s task behavior are not guaranteed to include all components involved in the model’s task behavior Mondorf et al. (2025). For details on the patching procedure, we refer to Appendix D.4.
@@
Figure 7: The linear probe’s accuracy across all layers of Qwen-2.5-1.5B at selected token positions.
@@ 4.3 Dissociation of Arithmetic Validation and Computation
Our findings presented in Section 4.2 suggest that the considered models tend to rely on surface-level consistency checks rather than re-evaluating the given arithmetic equation and comparing the result with the final numeric answer. Notably, we find that all models achieve 100% accuracy in predicting the correct result of the arithmetic equation provided in each prompt (e.g., $5+8=\text{?}$). However, we hypothesize that this correctly predicted result is not used for validation. To better understand the relationship between the models’ arithmetic computation and validation procedures, we identify another set of circuits responsible for computing the correct arithmetic result at the position of the arithmetic equation, following a similar procedure as described in Section 3. For further details on the identification process, please refer to Appendix D.3. Interestingly, we find that for all models, the identified arithmetic circuits predominantly contain edges in higher layers (due to space constraints, visualizations of these circuits are provided in Figures 22 to 25 in the Appendix). This structural dissociation between the circuits responsible for arithmetic computation and those involved in validation seems to explain the models’ difficulties in detecting basic arithmetic errors. Specifically, although the models can successfully re-compute the result of a given arithmetic equation, the final arithmetic outcome is not fully encoded when the model checks for numeric consistency in middle layers. To support this hypothesis, we train a linear probe to predict the correct arithmetic result based on the hidden states of the model’s residual stream (for training details, see Appendix D.5). Figure 7 shows the probe’s accuracy across different layers of Qwen-2.5-1.5B and selected token positions. Only in the higher layers (after the consistency check) does the model’s residual stream linearly encode information about the correct arithmetic result. Interestingly, similar patterns are observed for other models, too (see Figure 20 in the Appendix).
@@
Figure 8: Error detection accuracy of Qwen-2.5-1.5B on (clean, corrupt) prompt pairs where (left) the clean prompt contains a consistent error at both error positions, and (right) an error is present only at the position of the arithmetic result. The blue intervention bar denotes the result after adding the hidden representation of the residual stream in layer 22 (at [result-first]) to the residual stream of layer 1 (at [result-second]).
@@
We demonstrate that by “bridging” the gap between arithmetic computation and validation, the model’s error detection capacity for prompts with consistent errors at both error positions can be significantly enhanced. Figure 8 shows the error detection accuracy of Qwen-2.5-1.5B before and after we add the hidden representation of the residual stream from layer 22 at token position [result-first] to the residual stream of layer 1 at token position [result-second] (essentially “moving” information from the top yellow layers in Figure 7 to lower layers). Notably, this approach improves the model’s ability to detect consistent errors by 81%. Furthermore, accuracy on samples containing errors only at the position of the arithmetic result does not decline markedly, suggesting that the added representation enhances rather than overwrites existing information. A similar trend for other models is presented in Figure 21 in the Appendix.
Figure 9: Error detection accuracy of Qwen-2.5-1.5B before and after patching six consistency heads. We evaluate the model on consistent-error prompts related to subtraction (left) or division (right) and patch activations from a single-error run. As a control, we patch six randomly chosen attention heads.
4.4 Consistency Heads in Other Arithmetic Operations
Given the important role of consistency heads in our experiments on addition, we conduct further analyses to investigate whether their influence extends to other arithmetic operations. To this end, we generate new math word problems involving subtraction, multiplication, and division, using the same template structure with the two key error positions at the arithmetic result and the final numeric answer. [^6] See Table 10 in the Appendix for the full set of templates and variables considered when generating data for other arithmetic operations. As for addition, we first evaluate the models’ error detection accuracy on 1,000 prompt pairs per template. Similarly, we find that all models (except for Phi-3-Mini-4k-Instruct) encounter greater difficulty in detecting consistent errors at both error positions compared to single-position errors. Complete results for all models and operations are reported in Table 7 in the Appendix.
Next, for models that struggle with consistent errors, we replicate the consistency head patching experiment presented in Section 4.2, this time for prompts involving other arithmetic operations. Focusing on Qwen-2.5-1.5B, we evaluate the model on subtraction and division prompts with consistent errors at both positions, where it incorrectly judged the reasoning as “valid,” and patch the latent activations of the consistency heads with those obtained from prompts containing an error only at the arithmetic result. Notably, across all operations, we patch the same consistency heads initially identified with prompts involving additions (Table 4 in the Appendix). Figure 9 shows the change in error detection accuracy, indicating that the causal role of consistency heads in error detection extends beyond addition. Similar effects are observed for Qwen-2.5-Math-1.5B and Llama-3.2-3B (Figure 19 in the Appendix).
Finally, we find that some larger models, such as Llama-3.1-70B, similarly struggle with detecting consistent arithmetic errors, as shown in Table 8 in the Appendix. This may indicate that, to some extent, these models also rely on consistency heads—a direction that future work could explore.
@@ 5 Related Work
Self-correction in LLMs refers to the ability of models to correct their own generated output (Kamoi et al., 2024b; Huang et al., 2024; Madaan et al., 2023). Recent studies (Tyen et al., 2024; Kamoi et al., 2024a) suggest that LLMs tend to struggle with intrinsic self-correction, especially with detecting errors in their own output (Huang et al., 2024; Tyen et al., 2024; Kamoi et al., 2024b). While most studies focus on improving the models’ ability to self-correct (Kamoi et al., 2024b; Madaan et al., 2023; Chen et al., 2024; Zhao et al., 2023), we study error detection from a mechanistic point of view.
@@
The underlying processes of arithmetic reasoning and error detection have been studied independently in LLMs so far. Several studies (Stolfo et al., 2023; Zhang et al., 2024; Nikankin et al., 2024) use causal mediation analysis (Pearl, 2001) to identify circuits that account for how LLMs process arithmetic operations. As of now, only few studies have analyzed self-correction in LLMs beyond the models’ generated output (Li et al., 2024; Liu et al., 2024b).
@@ 7 Limitations
As discussed in Sections 1 and 3, our analysis is limited to four smaller-sized LLMs. Although we observe consistent patterns across various model architectures, sizes, and fine-tuning procedures (particularly within the mathematical domain), future research could investigate how these findings extend to larger models with more advanced arithmetic capabilities. Our behavioral experiments with models such as Llama-3.1-70B-Instruct and Qwen-2.5-32B-Instruct (see Table 8 in the Appendix) show that even larger LLMs tend to struggle more with consistent errors at both error positions, compared to detecting an error present only at the position of the arithmetic result or the final answer. This may indicate that, similarly to smaller models, these models—at least to some extent—rely on consistency heads that are susceptible to the validation gap. We believe this is a promising direction for future work to explore.
@@ Acknowledgments
We would like to thank ELLIS—the European Laboratory for Learning and Intelligent Systems—as the collaboration that led to this paper started at a workshop organized by Raquel Fernández and Sandro Pezzelle at MFO, the Oberwolfach Research Institute for Mathematics in the German Black Forest, on behalf of the ELLIS NLP program. We are also grateful to the members of the MaiNLP lab for their valuable feedback, with special thanks to Michael Hedderich, Robert Litschko, Diego Frassinelli, Rob van der Goot, Siyao Peng, Yang Janet Liu, Xinpeng Wang, Verena Blaschke, Raoyuan Zhao, Elena Senger, Felicia Körner, Soh-Eun Shim, Andreas Säuberli, Florian Eichin, Shijia Zhou, and Beiduo Chen. We further thank the anonymous reviewers for their insightful comments and suggestions. RB acknowledges the generous support of Amazon Alexa, whose donation helped make this work possible. Finally, we acknowledge the support for BP through the ERC Consolidator Grant DIALECT 101043235.
@@ B.3 Circuit Identification Process
Appendix C Dataset
C.1 Templates
We generate our dataset of clean and corrupt prompts based on the templates in Table 9. These templates have a set of variables, namely [instruction], [person], [object], [verb], [pronoun], [num1], [num2], [num3], each of which can be assigned different values. Table 11 lists all possible values. For the numerical variables, we use single-digit numerical values ([num1] and [num2]) that add to a two-digit arithmetic result ([num3]). To ensure that each variable assignment occupies the same position in the token sequence within a template, we retain only variables that are tokenized as a single token for each model. For the [instruction] variable, we include only instructions that share the same number of tokens. Finally, for the [correct_pair] variable, we select assignments where labels are tokenized as a single token across models.
For other operations (subtraction, multiplication, and division) we follow the same procedure and construct eight templates for each operation (see Table 10). These templates use the same variable structure as addition, with subtraction including also a [verb] variable. For the numerical variables, we restrict number pairs to ensure two-digit results: subtraction uses 2-digit minus 1-digit inputs yielding a 2-digit result, multiplication uses 2-digit times 1-digit inputs yielding a 2-digit result, and division uses 2-digit dividends and 1-digit divisors producing 2-digit quotients.
C.2 Aligning Token Positions Across Templates
Appendix D Experiment Details
Model Attention Heads
Consistency
Qwen-2.5-1.5B-Instruct L12H0, L12H2, L12H10, L13H0, L13H1, L13H10
Qwen-2.5-Math-1.5B-Instruct L13H0, L13H1
Llama-3.2-3B-Instruct L4H14, L5H3, L7H11, L8H1, L10H5, L10H18
Phi-3-Mini-4k-Instruct L10H1, L10H5, L10H14, L14H19, L16H18
Random
Qwen-2.5-1.5B-Instruct L1H0, L4H9, L9H6, L10H5, L11H8, L27H9
Qwen-2.5-Math-1.5B-Instruct L4H9, L27H5
Llama-3.2-3B-Instruct L1H1, L4H19, L9H16, L10H15, L11H23, L27H12
Phi-3-Mini-4k-Instruct L0H25, L3H18, L8H25, L20H19, L23H28